OpenStudy (anonymous):

write an equation in slope intercept form (-2,1) slope is 7

OpenStudy (anonymous):

there are two ways to do this.... (1) you can use point slope formula... y - y1 = m(x - x1) slope(m) = 7 (-2,1) x1 = -2 and y1 = 1 and just sub those in and put your equation in y = mx + b form. or (2) y = mx + b slope(m) = 7 (-2,1) x = -2 and y = 1 sub that in and solve for b which way do you choose to do it and I can help you ?

OpenStudy (anonymous):

I figured it out already, thank you tho!

OpenStudy (anonymous):

ok...good to hear :)